# TermSight: Making Service Contracts Approachable

### v1

@@ Appendix A Iterative Design
7 participants preferred and primarily used the condensed layout of the Summary Snippets during the reading session. They noted that the purpose of the Summary Snippets was to support navigation, and the condensed layout made it easier to gain an overview and identify relevant information in a policy. In contrast, the in-context layout required significantly more scrolling, which participants found overwhelming. As a result, in the final version of TermSight, we used the condensed layout for the Summary Snippets (Figure  1).
@@
All participants found the colors to be intuitive and effective in highlighting power and relevance, helping them decide what to read. However, 3 participants pointed out that while the power dimension was intuitive, the relevance dimension with 3 saturation levels was challenging to differentiate as there would be 9 colors in total. To reduce visual complexity, the final version of TermSight included only two saturation levels (High vs. Low) for relevance (Figure  2).
@@
Appendix B Additional Implementation Details of TermSight
In this section, we provide additional implementation details of TermSight.
@@ B.2. Obtaining Summary Snippets and Information Snippets
The prompt used to obtain the Summary Snippets and Information Snippets is detailed in Figure  21. We constrained the summary snippets to be short because prior works have found that adding short summaries (10-20 words) under search results was more effective for navigational and information-seeking tasks compared to longer summaries, which can be harder to skim  (Cutrell and Guan, 2007; Sweeney and Crestani, 2006). This length constraint was further optimized through prompt engineering. We noticed that setting summaries shorter than 12 words led to excessive fragmentation (i.e., too many summary snippets), while longer summaries tried to cover too much information, making them visually dense and less skimmable.
@@ B.3. Classifying Information Snippets
In TermSight, two classifications were performed for each Information Snippet using GPT-4o with few-shot prompting (Figure  22 and  23). For the classification of Power, each Information Snippet was classified by the degree of control or agency it grants to the Service Provider or the User (Categories: Service, Neutral, User). For the classification of Relevance, Information Snippets were classified based on their relevance to the user persona (Categories: High, Low). User personas used for the study are included in Figure  17 and  18.
@@ B.4. Phrase Scope
B.4.1. Identifying unfamiliar or vague phrases
For each document chunk, we employ two few-shot prompts to GPT-4o to identify potentially unfamiliar (Figure  24) or vague phrases (Figure  25). Few-shot examples for identifying unfamiliar phrases were selected based on phrases participants found challenging in our formative study, which included legal jargon, copyright licenses, and privacy-related terminology. Few-shot examples for identifying vague phrases were informed by vague phrases participants identified in the formative study, as well as prior works analyzing vagueness in Terms of Service (ToS) documents  (Lebanoff and Liu, 2018; Bhatia et al., 2016; Yerby and Vaughn, 2022). Both prompts were applied to each document chunk, producing two sets of identified phrases. Since a phrase may be both unfamiliar and vague, the union of these two sets was taken as the final set of identified phrases for a given chunk of text.
@@ B.4.2. Generating phrase definitions and answers to user questions
We used a retrieval-augmented question answering approach  (Fok et al., 2024a) to generate the phrase definitions. Figure  26 shows the prompt used to generate the definitions after retrieving potentially relevant document chunks. The prompt explicitly instructs GPT-4o to produce a concise in-context definition of the input phrase and output a list of the retrieved chunks referenced in generating the definition. In the case when the definition is generated without referencing any of the retrieved chunks, it is still retained. TermSight intends to provide plausible definitions even when explicit definitions are absent from the document. This design decision is also supported by prior research demonstrating the capabilities of large language models in generating term definitions that are comparable to traditional dictionaries  (Yin and Skiena, 2023; Jhirad et al., 2023). When users ask additional questions, the same retrieval-augmented question answering pipeline used for generating definitions is applied, with the only difference being the question asked (Figure  28). If the answer is not specified in the retrieved chunks, the system generates a plausible answer based on the phrase context, similar to the definitions.
@@ B.4.3. Generating scenarios
Our needfinding reveals the need to help readers not only interpret the definition of a phrase but to envision potential implications. As a result, we leveraged GPT-4o to generate customized scenarios of potential implications based on user persona. The prompt is specified in Figure  27.
@@ Appendix C Evaluation of TermSight Output
We manually evaluated TermSight’s feature pipelines and overall performance. Our goal was not to assess advances in system performance, but to verify that the system reliably produced meaningful outputs that could support ToS reading.
C.1. Evaluating the Classification of Information Snippets
We sampled 116 Information Snippets with stratified sampling from all the sub-policies for each service included in the user study. For each Summary Snippet, we evaluated the classification of power and relevance. Of the 116 Information Snippets, 27 were classified by TermSight as favoring the user, 22 were classified as favoring the service, and 67 were classified as neutral. In addition, 59 were classified as of high relevance to the provided user persona and 57 were classified to be of low relevance.
@@ C.1. Evaluating the Classification of Information Snippets
Our evaluation revealed 3 instances where the power classifications were imperfect, mainly caused by the lack of context in the input information snippet. For example, the snippet (‘All Buy Now purchases in a ServiceY Show are final and binding’) was classified by TermSight as favoring the service. In the generated explanation, the assumption was that the statement meant that the user can not return or refund the purchase. However, users can return and get a refund if there are problems with the purchase (e.g., an item doesn’t match its description) as stated in the service’s return policy, making this snippet more neutral. The return policy was not included in the snippet context, as a result, this neutral statement was considered more service-oriented.
@@
For the classification of relevance, there were 2 instances, out of 116, where irrelevant information snippets were classified as relevant to the persona. In both cases, the snippets were more relevant to sellers, not buyers, even though the persona given was that of a buyer. There were 11 instances where the information snippets classified as relevant were indirectly relevant to the input user personas. For example, sellers have to pay a fee to the platform after selling an item. Despite this information being more relevant to sellers, TermSight classified it as relevant to buyers as it exposed the potential hidden fees that sellers might include as part of their listing price.
C.2. Evaluating Term Definitions and Scenarios
We randomly sampled 113 generated definitions and scenarios using stratified sampling from all the sub-policies for each of the two ToS augmented for the user study. These were evaluated by members of the research team. Out of the 113 sampled phrases, 46 were identified by TermSight as vague phrases, 6 were identified as both jargon and vague phrases, and 61 were identified as jargon.
@@ C.2. Evaluating Term Definitions and Scenarios
Our evaluation revealed that all the generated definitions were correct, but out of the 113 definitions, 4 were only general definitions of the phrase not specific to the ToS. On further inspection, we identified that this was because these vague phrases were not explicitly defined anywhere in the ToS. Additionally, their meanings are service-specific and cannot be extrapolated from common sense or inferred by large language models (LLMs). For example, for the phrase ‘aggregated anonymized statistics’, TermSight provided a general definition of what it might mean to aggregate and anonymize user data. However, the specific details—such as what user data are being aggregated—were not specified in the ToS. Rather than a limitation of TermSight, these imperfections highlight the ill-defined nature of the language used in ToS.
@@
For the generated scenarios of phrases relevant to the user persona, we found one instance where the scenario was factually incorrect based on the input context. The input context and phrase stated that users do not gain ownership rights by downloading content from the service. However, the scenario claimed that users might lose ownership rights over the content they create by uploading it to the service. We also noticed that when the passages or phrases target a different audience (e.g., developers) than the user persona used to generate the scenarios (i.e., a lay user), the generated scenarios become less relevant or useful. We did not regenerate these imperfections to keep the user experience realistic to real-world settings where the LLM outputs are not guaranteed to be perfect. Participants in our user study were informed that AI output can be imperfect. This specific incorrect scenario was not accessed by any participants in the study.
Appendix D Bayesian Analysis Model Details
In this section, we provide additional details on the models used in the paper. All models used the DAG in Figure 9 to inform the model. Next, in § D.1, we describe the model used to estimate the effect of the treatment on the responses related to the experience of using the treatment interface. In § D.2, we describe the model used to estimate the effect of the treatment on the comprehension outcomes i.e., the quiz questions. Finally, in § D.3, we describe the model used to estimate the effect of the treatment on the experiment subjects’ ability to recall facts correctly.  Figure 19 and  Figure 20 show the forest plots of model coefficients for the experience and comprehension outcomes.
D.1. Modeling Experience Outcomes
We asked the experiment subjects to rate their experience of using the treatment or the baseline interface on a Likert scale from 1 to 5 (L=5 outcome values). There were N subjects and we asked each subject K (K=7) questions. Since the likert scale is ordinal, we modeled the experience outcomes using an ordinal regression model. The DAG in Figure 9 indicates that the experimental condition alone affects the experience outcomes. There were three randomizations: whether the subject was shown the treatment or control interface, and whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second. We modeled the experience outcomes using a cumulative logit model with a linear link function.
The model is given by:
[…]
In eq. 7, we further ensured that the samples are ordered i.e., $\kappa_{1}<\kappa_{2}<\dots<\kappa_{L-1}$. The priors are conservative. For example, on the logit scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse logistic to outcome probability range of [0.04, 0.95].
D.2. Modeling Comprehension Outcomes
Like earlier, there were N subjects, and we asked each subject K (K=6) questions to test their comprehension of the ToS. Thus, the outcome is a binary variable with 1 indicating a correct answer and 0 indicating an incorrect answer. The DAG in Figure 9 indicates that the experimental condition alone affects the comprehension outcomes. There were three randomizations: whether the subject was shown the treatment or control interface, whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second. We modeled the comprehension outcomes using a logistic regression model. Notice that we created a joint variable $c_{j,k,l}$ for the coefficients corresponding to the question, service, and treatment/control. We did this because the questions were different across the two services used in the experiment. We included the treatment/control variable to be able to easily estimate the effect of the treatment on the comprehension outcomes per question.
The model is given by:
[…]
As in the previous model, the priors are conservative. For example, on the logit scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse logistic to outcome probability range of [0.04, 0.95].
D.3. Modeling Recall Outcomes
There were N subjects, and we asked each subject a question on recall facts from the ToS. The outcome that we measure is the number of correctly recalled facts, making the outcome a count variable. Since many of the subjects could not recall any facts correctly, we modeled the recall outcomes using a Zero Inflated Poisson regression model. The DAG in Figure 9 indicates that the experimental condition alone affects the comprehension outcomes. As before, there were three randomizations: whether the subject was shown the treatment or control interface, whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second.
The model is given by:
[…]
As in the previous model, the priors are conservative. For example, on the Log scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse log to outcome range of [0.05, 21]. Notice that we model the number of correctly recalled facts, and thus is a conservative prior. Given that most subjects could not recall any facts, the zero inflation parameter is set to a weakly informative prior using a Beta distribution.
Appendix E Study Materials: Formative Study
@@
  • What information were you interested in in ToS? Both from your prior experience in reading ToS or this ToS?
  • Imagine if you have a magic wand that can transform the Terms of Service in whatever ways you want. How would you transform the Terms of Service?
Appendix F Study Materials: User Study
F.1. User Persona
Two personas were given to the participants during the user study for the social media service (Figure  17) and e-commerce service (Figure  18). The same personas were used for features of TermSight to classify relevance and generate personalized scenarios. The personas were designed based on information participants in the formative study described caring about.
User Persona for Social Media site: Content consumer who posts personal content Imagine you are a lay user of social media platforms. You are over 18 years old and located in the United States. Your usage of Social Media sites: You spend most of your time on the platform scrolling through feeds, liking posts, chatting with other users, and sharing personal content such as photos. Things you care about when using Social Media Sites: You care about Privacy, particularly what data are being collected and how your data can be used and shared. You care about what the service can do with user-generated content such as licenses over user content or advertising with user content. You care about potential liabilities when using ServiceX.
Figure 17. User Persona for Social Media site: Content consumer who posts personal content
User Persona for E-commerce site: Buyer who rarely posts reviews Imagine you are a lay user of E-commerce platforms. You are over 18 years old and located in the United States. Your usage of E-commerce sites: You typically engage with the E-commerce platform to buy new or used items from other users. You rarely post any reviews or content on the service. Things you care about when using E-commerce sites: You care about information related to making purchases, refunds, returns, user protection policies, termination, arbitration, and liabilities. You also care about Privacy, particularly what data are being collected and how your data can be used and shared.
Figure 18. User Persona for E-commerce site: Buyer who rarely posts reviews
F.2. Pre-survey Questions
Questions that were asked only once at the beginning of the interview:
  • For how many online platforms have you read their Terms and Conditions (T&C) before? [None (0), Few (1-3), Some (4-6), Many (7-9), A lot (¿10)]
  • How familiar are you with Terms and Conditions (T&C) for online platforms? (5-point likert rating)
@@ F.2. Pre-survey Questions
F.3. 5-point Likert Ratings of Reading Experience
@@
  • E1: I’m interested in spending more time reading the service’s Terms and Conditions (T&C) with the current interface and wish to get a link after the study.
  • E2: How hard did you have to work to read the T&C?
@@
  • E5: How confident are you that you got all the relevant information from the T&C (including sub-policies)?
  • E6: How much do you feel like you understood the T&C?
  • E7: How much would you be willing to read the T&C of other services with this interface?
F.4. Semi-structured Interview Questions
Below are questions that were asked after each of the two reading sessions, except the last question which was only asked after the second reading session where participants would compare and contrast both interfaces.
  • Describe your experience reading the Terms of Service using the interface.
@@ F.4. Semi-structured Interview Questions
Question Correct Answer Relevant Passage

### v2

@@ Appendix A Iterative Design Study
7 participants preferred and primarily used the condensed layout of the Summary Snippets during the reading session. They noted that the purpose of the Summary Snippets was to support navigation, and the condensed layout made it easier to gain an overview and identify relevant information in a policy. In contrast, the in-context layout required significantly more scrolling, which participants found overwhelming. As a result, in the final version of TermSight, we used the condensed layout (Figure  1).
@@
All participants found the colors to be intuitive and effective in highlighting power and relevance, helping them decide what to read. However, 3 participants pointed out that while the power dimension was intuitive, the relevance dimension with 3 saturation levels was challenging to differentiate, as there would be 9 colors in total. To reduce visual complexity, the final version of TermSight included only two saturation levels (High vs. Low) for relevance (Figure  2).
@@
Appendix B Additional Implementation Details
@@ B.2. Obtaining Summary Snippets and Information Snippets
The prompt used to obtain the Summary Snippets and Information Snippets is detailed in Figure  17. We constrained the Summary Snippets to be short because prior works have found that adding short summaries (10-20 words) under search results was more effective for navigational and information-seeking tasks compared to longer summaries, which can be harder to skim  (Cutrell and Guan, 2007; Sweeney and Crestani, 2006).
@@ B.3. Classifying Information Snippets
In TermSight, two classifications were performed for each Information Snippet using GPT-4o with few-shot prompting (Figure  18 and  19). For the classification of Power, each Information Snippet was classified by the degree of control or agency it grants to the Service Provider or the User (Categories: Service, Neutral, User). For the classification of Relevance, Information Snippets were classified based on their relevance to the user persona (Categories: High, Low). User personas used for the study are included in Appendix  E.2.
@@ B.4. Phrase Scope
B.4.1. Identifying unfamiliar or ambiguous phrases
For each document chunk, we employ two few-shot prompts to GPT-4o to identify potentially unfamiliar (Figure  20) or vague phrases (Figure  21). Few-shot examples for identifying unfamiliar and ambiguous phrases were selected based on phrases participants found challenging in our formative study. Both prompts were applied to each document chunk, producing two sets of identified phrases. Since a phrase may be both unfamiliar and vague, the union of these two sets was taken as the final set of identified phrases for a given chunk of text.
@@ B.4.2. Generating phrase definitions and answers to user questions
We first retrieve potentially relevant document chunks from the vector database. The database query is framed as a question: “What does {input phrase} refer to in the sentence: {phrase context}. Here, phrase context refers to the chunk of text containing the input phrase. Then, the query is embedded using OpenAI’s text-embedding-3-small model, and the top 15 chunks from the vector database are retrieved based on cosine similarity. Figure  22 shows the prompt used to generate the definitions after retrieving potentially relevant document chunks. When users ask additional questions, the same retrieval-augmented question answering pipeline used for generating definitions is applied, with the only difference being the question asked (Figure  24).
@@ B.4.3. Generating scenarios
We leveraged GPT-4o to generate customized scenarios of potential implications based on user persona. The prompt is specified in Figure  23.
Classification Input Snippet Output Imperfection
Power “All Buy Now purchases in a ServiceY Show are final and binding.” Service Lack of enough context in the input. Users do have the option to return. It’s a neutral clause.
Power “Use public content for any illegal, deceptive, unethical, false, misleading, or improper purpose, including the infringement of third-party intellectual property rights.” Neutral Lack of enough context in the input. The clause describes what third party licensee cannot do with user content. It’s a user-benefiting clause.
Relevance “Your earnings are based on the listing price and actual earnings will vary based on the final order price, Seller discounts, and any other applicable taxes and discounts.” High Lack of enough context in the input. This clause is for sellers and is less relevant to the input buyer persona.
Relevance “ServiceY cannot guarantee that a ServiceY consignment listing will be sold or that a certain sales amount will be earned for individual items or an entire shipment.” High Lack of enough context in the input. This clause is for sellers and is less relevant to the input buyer persona.
Table 1. Examples of imperfect classifications of Power and Relevance.
Examples of imperfect classifications of Power and Relevance. The four columns describe the type of classification (Relevance or Power), model input, model output, and the imperfection.
@@ Appendix C Evaluation of TermSight Output
C.1. Classification of Information Snippets
@@
Out of the sampled 116 snippets, our evaluation revealed 3 instances where the power classifications were imperfect, mainly caused by the lack of context in the input Information Snippet. For example, the snippet (‘All Buy Now purchases in a ServiceY Show are final and binding’) was classified by TermSight as favoring the service. In the generated explanation, the assumption was that the statement meant that the user can not return or refund the purchase. However, users can return and get a refund if there are problems with the purchase (e.g., an item doesn’t match its description) as stated in the service’s return policy, making this snippet more neutral. The return policy was not included in the snippet context; as a result, this neutral statement was considered more service-oriented. We didn’t observe any clause favoring the service provider being misclassified as neutral or favoring the user.
@@
For the classification of relevance, there were 2 instances, out of 116, where irrelevant Information Snippets were classified as relevant to the persona. In both cases, the snippets were more relevant to sellers, not buyers, even though the persona given was that of a buyer. There were 11 instances where the Information Snippets classified as relevant were indirectly relevant to the input user personas. For example, sellers have to pay a fee to the platform after selling an item. Despite this information being more relevant to sellers, TermSight classified it as relevant to buyers, as it exposed the potential hidden fees that sellers might include as part of their listing price. We show examples of imperfect classification of Power and Relevance in Table  1.
C.2. Term Definitions and Scenarios
@@
Our evaluation revealed that all the generated definitions were correct, but out of the 113 definitions, 4 were only general definitions of the phrase, not specific to the ToS. On further inspection, we identified that this was because these vague phrases were not explicitly defined anywhere in the ToS. Additionally, their meanings are service-specific and cannot be extrapolated from common sense or inferred by large language models (LLMs). For example, for the phrase ‘aggregated anonymized statistics’, TermSight provided a general definition of what it might mean to aggregate and anonymize user data. However, the specific details—such as what user data is being aggregated—were not specified in the ToS. Rather than a limitation of TermSight, these imperfections highlight the ill-defined nature of the language used in ToS.
@@
For the generated scenarios of phrases relevant to the user persona, we found one instance where the scenario was factually incorrect based on the input context. The input context and phrase stated that users do not gain ownership rights by downloading content from the service. However, the scenario claimed that users might lose ownership rights over the content they create by uploading it to the service. We also noticed that when the passages or phrases target a different audience (e.g., developers) than the user persona used to generate the scenarios (i.e., a lay user), the generated scenarios become less relevant or useful. We did not regenerate these imperfections to keep the user experience realistic in real-world settings where the LLM outputs are not guaranteed to be perfect. Participants in our user study were informed that AI output can be imperfect. This specific incorrect scenario was not accessed by any participants in the study.
Appendix D Study Materials: Formative Study
@@
  • What information were you interested in in ToS? Both from your prior experience in reading ToS and this ToS?
  • Imagine if you have a magic wand that can transform the ToS in whatever ways you want. How would you transform the Terms of Service?
Appendix E Study Materials: User Study
E.1. Baseline Interface
The baseline ToS reading interface used in the user study can be found in Figure  25.
E.2. User Persona
Two personas were given to the participants during the user study for the social media service (content consumer who posts personal content) and the e-commerce service (buyer who rarely posts reviews). The same personas were used for features of TermSight to classify relevance and generate personalized scenarios. The personas were designed based on information that participants in the formative study described caring about.
Persona: Content consumer who posts personal content Imagine you are a lay user of social media platforms. You are over 18 years old and located in the United States. Your usage of Social Media sites: You spend most of your time on the platform scrolling through feeds, liking posts, chatting with other users, and sharing personal content such as photos. Things you care about when using Social Media Sites: You care about Privacy, particularly what data is being collected and how your data can be used and shared. You care about what the service can do with user-generated content, such as licenses over user content or advertising with user content. You care about potential liabilities when using ServiceX.
Persona: Buyer who rarely posts reviews Imagine you are a lay user of E-commerce platforms. You are over 18 years old and located in the United States. Your usage of E-commerce sites: You typically engage with the E-commerce platform to buy new or used items from other users. You rarely post any reviews or content on the service. Things you care about when using E-commerce sites: You care about information related to making purchases, refunds, returns, user protection policies, termination, arbitration, and liabilities. You also care about Privacy, particularly what data is being collected and how your data can be used and shared.
E.3. Participant Demographics
Of the 20 participants, 13 self-identify as female and 7 as male. 1 participant had a high school degree, 3 had an associate degree, 12 had a bachelor’s degree, and 4 had a master’s degree. 4 participants were between the ages of 18-25, 5 were between 26-35, 5 were between 36-45, and 6 were between 46-55. When asked about how many ToS they’ve read, 4 participants read none, 5 read between 1–3, 5 read between 4–6, 1 read between 7–9, and 5 read greater than 10. Participants in the user study also found their given user persona for both services to highly align with their personal usage of the service and their personal value (Figure  13).
E.4. Pre-survey Questions
Questions that were asked once at the beginning of the interview:
  • For how many online platforms have you read their Terms of Service (ToS) before? [None (0), Few (1-3), Some (4-6), Many (7-9), A lot (¿10)]
  • How familiar are you with Terms of Service (ToS) for online platforms? (5-point Likert rating)
@@ E.4. Pre-survey Questions
Figure 13. Participants’ self-ratings of their familiarity with social media and e-commerce platforms and their alignment with the user persona given for each service. Most participants were familiar with both social media and e-commerce platforms and found the given persona to align with their personal usage of the service and personal value.
Participants’ self-ratings of their general familiarity with social media and e-commerce platforms and their alignment with the user persona given for each service. Most participants were familiar with both social media and e-commerce platforms in general and found the given persona to align with their personal usage of the service and personal value.
E.5. 5-point Likert Ratings of Reading Experience
@@
  • E1: I’m interested in spending more time reading the service’s Terms of Service (ToS) with the current interface and wish to get a link after the study.
  • E2: How hard did you have to work to read the ToS?
@@
  • E5: How confident are you that you got all the relevant information from the ToS (including sub-policies)?
  • E6: How much do you feel like you understood the ToS?
  • E7: How much would you be willing to read the ToS of other services with this interface?
E.6. Semi-structured Interview Questions
Below are questions that were asked after each of the two reading sessions, except the last question. The last question was only asked after the second reading session, where participants would compare and contrast both interfaces.
  • Describe your experience reading the ToS using the interface.
@@ E.6. Semi-structured Interview Questions
Appendix F Bayesian Analysis Model Details and Additional Analysis
In this section, we provide details on the Bayesian models. We first explain the concepts of structural causal framework and DAG in §F.1. All models were informed by the DAG in Figure  6. Next, in §F.2, we describe the model used to estimate the effect of the treatment on the user experience. In §F.3, we describe the model used to estimate the effect of the treatment on the comprehension outcomes, i.e., the quiz questions. Finally, in §F.4, we describe the model used to estimate the effect of the treatment on the experiment subjects’ ability to recall facts correctly. Figure  14 and  15 show the forest plots of model coefficients for the experience and comprehension outcomes. Figure  16 presents forest plots illustrating the effect of ordering on user experience, comprehension, and recall.
F.1. Structural Causal Framework
We used a structural causal framework popularized by Pearl (2009) to better understand the effects of the treatment. A structural causal model framework comprises exogenous variables $U$, endogenous variables $V$, and a set of functions $F$ that represent the physical process that results in the values of the endogenous variables. For example, the question $v_{i}=f(v,u)$ implies that the values of $v$ and $u$ completely determine the value of $v_{i}$ through the function $f$. The structural causal model can be represented as a Directed Acyclic Graph (DAG) where the nodes represent the variables and the edges represent the causal relationships among the variables. Thus, in the preceding example, $v$ and $u$ are the parents of $v_{i}$ in the DAG. The structural causal model framework is non-parametric, but frequently used with generalized linear models to estimate the functions $f$. The structural causal framework has connections with structural equation models (Wright, 1934) used in the Economic Sciences, though the latter were developed with linear relationships in mind. Figure  6 shows the DAG for the study.
F.2. Modeling Experience Outcomes
We asked the experiment subjects to rate their experience of using the treatment or the baseline interface on a Likert scale from 1 to 5 (L=5 outcome values). There were N subjects, and we asked each subject K (K=7) questions. Since the Likert scale is ordinal, we modeled the experience outcomes using an ordinal regression model. The DAG in Figure  6 indicates that the experimental condition alone affects the experience outcomes. There were three randomizations: whether the subject was shown the treatment or control interface (i.e., interface type), whether the subject was shown the social media ToS or the e-commerce ToS (i.e., service type), and whether the treatment was shown first or second (i.e., presentation order). We modeled the experience outcomes using a cumulative logit model with a linear link function:
(1) $Y_{i,j,k,l,m}$ $\sim\mathrm{OrderedLogistic}(\kappa,\phi_{i,j,k,l,m})\quad$
(2) $\phi_{i,j,k,l,m}$ $=\alpha_{i}+c_{j,k}+s_{l}+o_{m}\quad$
Below, we specify the priors for $\alpha_{i}$ (intercept, one per participant), $c_{j,k}$ (joint variable for the interface type and user experience item), $s_{l}$ (service type), $o_{m}$ (presentation order), and $\kappa_{i}$ (cutpoints).
(3) $\alpha_{i}$ $\sim\mathrm{Normal}(0,1),i\in\{1,\dots,N\}\quad$
(4) $c_{j,k}$ $\sim\mathrm{Normal}(0,1),j\in\{1,2\},k\in\{1,\dots,K\}\quad$
(5) $s_{l}$ $\sim\mathrm{Normal}(0,1),l\in\{1,2\}\quad$
(6) $o_{m}$ $\sim\mathrm{Normal}(0,1),m\in\{1,2\}\quad$
(7) $\kappa_{i}$ $\sim\mathrm{Normal}(0,1),i\in\{1,\dots,L-1\}\quad$
In eq (7), we further ensured that the samples are ordered, i.e., $\kappa_{1}<\kappa_{2}<\dots<\kappa_{L-1}$. The priors are conservative. For example, on the logit scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse logistic to the outcome probability range of [0.04, 0.95].
F.3. Modeling Comprehension Outcomes
There were N subjects, and we asked each subject K (K=6) questions to test their comprehension. The outcome is a binary variable with 1 indicating a correct answer and 0 indicating an incorrect answer. The DAG in Figure  6 indicates that the experimental condition alone affects the comprehension outcomes. There were three randomizations: whether the subject was shown the treatment or control interface, whether the subject was shown the social media ToS or the e-commerce ToS, and whether the treatment was shown first or second. We modeled the comprehension outcomes using a logistic regression model. Notice that we created a joint variable $c_{j,k,l}$ for the coefficients corresponding to the interface type, question, and service type. We did this because the comprehension questions were different across the two services used in the experiment. We included the interface variable (i.e., treatment vs. control) to be able to easily estimate the effect of the treatment on the comprehension outcomes per question. The model is given by:
(8) $Y_{i,j,k,l,m}$ $\sim\mathrm{Binomial}(p_{i,j,k,l,m})\quad$
(9) $\mathrm{Logistic}(p_{i,j,k,l,m})$ $=\alpha_{i}+c_{j,k,l}+o_{m}\quad$
Below, we specify the priors for $\alpha_{i}$ (intercept, one per participant), $c_{j,k,l}$ (joint variable for the interface type, comprehension question, and service type), and $o_{m}$ (presentation order).
(10) $\alpha_{i}$ $\sim\mathrm{Normal}(0,1),i\in\{1,\dots,N\}\quad$
(11) $c_{j,k,l}$ $\sim\mathrm{Normal}(0,1),j\in\{1,2\},k\in\{1,\dots,K\},l\in\{1,2\}\quad$
(12) $o_{m}$ $\sim\mathrm{Normal}(0,1),m\in\{1,2\}\quad$
As in the previous model, the priors are conservative. For example, on the logit scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse logistic to the outcome probability range of [0.04, 0.95].
F.4. Modeling Recall Outcomes
There were N subjects, and we asked each subject to recall facts from the ToS. The outcome that we measure is the number of correctly recalled facts, making the outcome a count variable. Since many of the subjects could not recall any facts correctly, we modeled the recall outcomes using a Zero Inflated Poisson regression model. The DAG in Figure  6 indicates that the experimental condition alone affects the comprehension outcomes. As before, there were three randomizations: whether the subject was shown the treatment or control interface, whether the subject was shown the social media ToS or the shopping ToS, and whether the treatment was shown first or second. The model is given by:
(13) $Y_{i,j,k,l}$ $\sim\mathrm{ZeroInflatedPoisson}(\lambda_{i,j,k,l},\phi)\quad$
(14) $\mathrm{Log}(\lambda_{i,j,k,l})$ $=\alpha_{i}+c_{j}+s_{k}+o_{l}\quad$
Below, we specify the priors for $\alpha_{i}$ (intercept, one per participant), $c_{j}$ (interface type), $s_{k}$ (service type), $o_{l}$ (presentation order), and $\phi$ (probability of zero inflation).
(15) $\alpha_{i}$ $\sim\mathrm{Normal}(0,1),i\in\{1,\dots,N\}\quad$
(16) $c_{j}$ $\sim\mathrm{Normal}(0,1),j\in\{1,2\}\quad$
(17) $s_{k}$ $\sim\mathrm{Normal}(0,1),k\in\{1,2\}\quad$
(18) $o_{l}$ $\sim\mathrm{Normal}(0,1),l\in\{1,2\}\quad$
(19) $\phi$ $\sim\mathrm{Beta}(2,2)\quad$
As in the previous model, the priors are conservative. For example, on the Log scale, a coefficient that is normally distributed with mean 0 and standard deviation 1, the range [-3, 3] covers 99% of the distribution and evaluates using inverse log to the outcome range of [0.05, 21]. Notice that we model the number of correctly recalled facts, and thus is a conservative prior. Given that most subjects could not recall any facts, the zero inflation parameter is set to a weakly informative prior using a Beta distribution.
Appendix G Supplemental Analysis for Comprehension Outcomes
We aimed to design the comprehension questions to have a single answer that best matches the clauses in the original text. However, we acknowledge that legal problems and contractual clauses may leave room for interpretation  (Marotta-Wurgler and Stein, 2025). As a result, we run an additional, more conservative analysis by removing any question that may contain room for interpretation on whether the original text would apply to the situation being asked in the question, leading to different answer choices. We removed 2 out of 6 questions for each service type (Social Media: Q1, Q4; Shopping: Q2, Q3). For all other questions, there is a single correct answer choice directly matching the original text, while all the other options either conflict with or were never mentioned in the original text. Below is one example of the questions removed from this analysis:
Social Media Q1: How can ServiceX use photographs you post for advertisements or promotions?
Original Answer: ServiceX can use photos you post in ads without your permission and is not obligated to attribute you as the creator.